OpenStudy (anonymous):

2cos^2theta=2+sin theta Please help me! Thank you so much, this is really appreciated, I'm having difficulty with these examples. :)

OpenStudy (anonymous):

\[\cos^2(\theta) = 1 - \sin^2(\theta)\]

OpenStudy (anonymous):

substitute that and solve like a quadratic

OpenStudy (mathmale):

Hello, Avi, Weren't there instructions with \[2(\cos^2\theta)=2+\sin \theta\] ? Please provide as much info as you can about what you want to accomplish. I can guess (but wish I didn't have to): You are to solve the given equation for theta: There is a very useful identity for (cos theta)^2. Mind doing some review, to find it? Get rid of the (cos theta)^2 and replace it with its equivalent.

OpenStudy (anonymous):

okay, thanks! I couldn't find my formula sheet, it didn't click at first that that was what I ought to do.

OpenStudy (anonymous):

and it was just shown as the problem, that was all the info I had, sorry.

OpenStudy (anonymous):

so, \[2(1-\sin^2(\theta)) = 2 + \sin(\theta)\]\[2 - 2\sin^2(\theta) = 2 + \sin(\theta)\]\[-2\sin^2(\theta)-\sin(\theta)=0\]\[(\sin(\theta))(-2\sin(\theta)-1)=0\] so \[\sin(\theta) = 0\]or\[\sin(\theta)=-1/2\]

OpenStudy (anonymous):

Thanks! The step by step was very informative!

OpenStudy (anonymous):

No Problem